OpenStudy (faman39):

A cube has a surface area of 16 square yards. What is the length of a side? How it can be 5 yrd ?

OpenStudy (jkristia):

I get \[x \approx 1.6329\]

OpenStudy (jkristia):

\[A = 6(length^2)\]

OpenStudy (faman39):

a. 5yd b. 12.5 yd c. 25 yd d. 37.3 yd

OpenStudy (faman39):

correct one 5yd but dont know how

OpenStudy (jkristia):

that is strange, and you are sure it is A=16, not 150? (which is 5)

OpenStudy (apoorvk):

how can it ever be 5yds? surface area of cube of edge 'a' is - (no. of faces) x (surface area of each face) =6*(a^2)

OpenStudy (faman39):

sorry it was 150

OpenStudy (faman39):

i wrote that wrong, sorry abt that

OpenStudy (apoorvk):

so, if side edge is 'a' \[16 = 6a^2\] \[or, a^2 = \frac {16}{6} = \frac 8 3 \] \[or, a = \sqrt {\frac 8 3 }\]

OpenStudy (apoorvk):

You mean '150' instead of '16'?

OpenStudy (faman39):

so we dont have to divide 8/3

OpenStudy (apoorvk):

Nope, then we have some thing like this: so, if side edge is 'a' \['50' =6a^2\] \[or,a^2=\frac {150} 6=25\] \[or,a=\sqrt{25}\] \[or, a = 5\] Pretty correct, I guess?

OpenStudy (apoorvk):

sorry, the first line would read '150' and NOT '50' - typo.

OpenStudy (faman39):

:)

OpenStudy (apoorvk):

:))